Evaluate the following integrals



(1) integral of {(10x^9+40x^4+3)√(x^10+8x^5+3x+5)} DX



(2) integral of {1/[(x-5)(x^2+4)]} dx

1
Expert's answer
2022-03-22T04:38:48-0400

1.

"\int\left(10 x^{9}+40 x^{4}+3\right) \sqrt{x^{10}+8 x^{5}+3 x+5} \mathrm{~d} x\\[4mm]
\text{ Substitute } u=x^{10}+8 x^{5}+3 x+5 \longrightarrow \frac{\mathrm{d} u}{\mathrm{~d} x}=10 x^{9}+40 x^{4}+3 \longrightarrow \mathrm{d} x=\frac{1}{10 x^{9}+40 x^{4}+3} \mathrm{~d} u:\\[4mm]
=\int \sqrt{u} \mathrm{~d} u\\[4mm]

\text{ Apply power rule:}\\[2mm]
\int u^{\mathrm{n}} \mathrm{d} u=\frac{u^{\mathbf{n}+1}}{\mathbf{n}+1} +C\text{ with }\mathbf{n}=\frac{1}{2} =\frac{2 u^{\frac{3}{2}}}{3} \text{ Substitute back } u=x^{10}+8 x^{5}+3 x+5 : \\[3mm] =\frac{2\left(x^{10}+8 x^{5}+3 x+5\right)^{\frac{3}{2}}}{3} +C\\[3mm]
\text{ Hence, }\\[3mm]
\begin{gathered}
\int\left(10 x^{9}+40 x^{4}+3\right) \sqrt{x^{10}+8 x^{5}+3 x+5} d x \\
=\frac{2\left(x^{10}+8 x^{5}+3 x+5\right)^{\frac{3}{2}}}{3}+C
\end{gathered}"


2.

"\int \frac{1}{(x-5)\left(x^{2}+4\right)} d x\\[4mm]
\text{ Perform partial fraction decomposition:}\\[2mm]
=\int\left(\frac{1}{29(x-5)}-\frac{x+5}{29\left(x^{2}+4\right)}\right) d x\\[2mm]
\text{ Apply linearity: } \\=\frac{1}{29} \int \frac{1}{x-5} \mathrm{~d} x-\frac{1}{29} \int \frac{x+5}{x^{2}+4} \mathrm{~d} x \\
\text{Now solving: }\\
\int \frac{1}{x-5} \mathrm{~d} x \\[4mm]
\frac{\mathrm{d} u}{\mathrm{~d} x}=1 \longrightarrow \mathrm{d} x=\mathrm{d} u =\int \frac{1}{u} \mathrm{~d} u\\[2mm]
\begin{gathered}
=\frac{1}{29} \int \frac{1}{x-5} \mathrm{~d} x-\frac{1}{29} \int \frac{x+5}{x^{2}+4} \mathrm{~d} x \\[3mm]
\text { Now solving: } \\[2mm]
\int \frac{1}{x-5} \mathrm{~d} x \\
\text { Substitute } u=x-5 \longrightarrow \frac{\mathrm{d} u}{\mathrm{~d} x}=1 \longrightarrow \mathrm{d} x=\mathrm{d} u:
\end{gathered}\\[2mm]
=\int \frac{1}{u} \mathrm{~d} u\\[3mm]
\text{ This is a standard integral:}\\[2mm]
=\ln (u)+C\\[3mm]
\text{ Substitute back} \,u = x-5\\
=\ln (x-5)+C\\[3mm]
\text{ Now solving:}\\[2mm]
\int \frac{x+5}{x^{2}+4} d x\\[3mm]
\text{ Expand:}\\[2mm]
=\int\left(\frac{x}{x^{2}+4}+\frac{5}{x^{2}+4}\right) d x\\
=\int \frac{x}{x^{2}+4} \mathrm{~d} x+5 \int \frac{1}{x^{2}+4} \mathrm{~d} x"

"\text{ Now solving:}\\[2mm]
\int \frac{x}{x^{2}+4} d x\\[3mm]
\text{ Substitute } u=x^{2}+4 \longrightarrow \frac{\mathrm{d} u}{\mathrm{~d} x}=2 x \longrightarrow \mathrm{d} x=\frac{1}{2 x} \mathrm{~d} u :\\[2mm]
=\frac{1}{2} \int \frac{1}{u} \mathrm{~d} u\\[3mm]
\text{ Now solving:}\\[2mm]
\int \frac{1}{u} \mathrm{~d} u\\[2mm]
\text{ Use previous result:}\\[2mm]
=\ln (u)+C_1\\[2mm]
\text{ Plug in solved integrals:}\\[2mm]
\begin{aligned}
&\frac{1}{2} \int \frac{1}{u} \mathrm{~d} u \\
&=\frac{\ln (u)}{2}+C
\end{aligned}
\\[2mm]
\text{Substitute back } u=x^{2}+4 :\\[2mm]
=\frac{\ln \left(x^{2}+4\right)}{2}+C\\[3mm]
\text{ Now solving:}\\[2mm]
\int \frac{1}{x^{2}+4} \mathrm{~d} x\\[2mm]
\text{ Substitute } u=\frac{x}{2} \longrightarrow \frac{\mathrm{d} u}{\mathrm{~d} x}=\frac{1}{2} \longrightarrow \mathrm{d} x=2 \mathrm{~d} u :\\[2mm]
=\int \frac{2}{4 u^{2}+4} d u"

"\text{ Simplify: }\\[2mm]
=\frac{1}{2} \int \frac{1}{u^{2}+1} \mathrm{~d} u\\[2mm]
\text{ Now solving:}\\[2mm]
\int \frac{1}{u^{2}+1} d u\\[2mm]
\text{ This is a standard integral:}
=\arctan (u)+C_1\\[2mm]
\text{ Plug in solved integrals:}\\[2mm]
\begin{gathered}
\frac{1}{2} \int \frac{1}{u^{2}+1} \mathrm{~d} u \\
=\frac{\arctan (u)}{2} +C\\
\text { Substitute back } u=\frac{x}{2} \text { : } \\
=\frac{\arctan \left(\frac{x}{2}\right)}{2}+C
\end{gathered}
\\[2mm]
\text{ Plug in solved integrals:}\\
\begin{aligned}
&\int \frac{x}{x^{2}+4} \mathrm{~d} x+5 \int \frac{1}{x^{2}+4} \mathrm{~d} x \\
&=\frac{\ln \left(x^{2}+4\right)}{2}+\frac{5 \arctan \left(\frac{x}{2}\right)}{2}+C
\end{aligned}\\[2mm]
\text{ Plug in solved integrals:}\\[2mm]
\begin{gathered}
\frac{1}{29} \int \frac{1}{x-5} \mathrm{~d} x-\frac{1}{29} \int \frac{x+5}{x^{2}+4} \mathrm{~d} x \\
=-\frac{\ln \left(x^{2}+4\right)}{58}-\frac{5 \arctan \left(\frac{x}{2}\right)}{58}+\frac{\ln (x-5)}{29}+C
\end{gathered}"

"\text{The problem is solved. Apply the absolute value function to arguments of}\\ \text{logarithm functions in order to extend the antiderivative's domain:}\\[2mm]
\begin{aligned}
& \int \frac{1}{(x-5)\left(x^{2}+4\right)} d x \\
=-& \frac{\ln \left(x^{2}+4\right)}{58}-\frac{5 \arctan \left(\frac{x}{2}\right)}{58}+\frac{\ln (|x-5|)}{29}+C \\
=&-\frac{\ln \left(x^{2}+4\right)+5 \arctan \left(\frac{x}{2}\right)-2 \ln (|x-5|)}{58}+C
\end{aligned}"